# Towards the Pedagogical Steering of Large Language Models for Tutoring: A Case Study with Modeling Productive Failure

### v1

@@ 5.3 Qualitative Insights from Students Feedback
Our field test demonstrated the ability of StratL tutor LLM to follow Productive Failure. We highlight in this section some insights gathered from our students’ feedback.
Robustness
@@
As discussed in Section 5.1, the Student State Tracing procedure can make labeling mistakes. These mistakes can then compromise the Intent Selection and result in the selection of contradictory intents. This can also happen if the Intent Selection is poorly designed. This is exemplified by the feedback of field study students: “The AI keeps on asking me to check my answer although it was correct”.
Grounding constraints
@@
A student who was assigned to the baseline with no intents (V2) complained that the tutor would not let them use a solution entirely different from the expected one: “The AI only listens to key points that it wants, and ignores everything else (e.g. my idea of using […])” A good tutor should encourage the use of an original method, but the model considers it a mistake. This highlights a real pedagogical problem, as in this case, the tutor is harmful to the learning by restricting the student’s exploration. This insight shows that Pedagogical Steering is not just about telling or not the answer, but rather about a collection of different pedagogical properties LLMs lack.
6 Conclusion and Limitations
@@
In this work, we identify the pedagogical limitations of LLMs in multi-turn conversational tutoring and present it as the Pedagogical Steering problem. We introduce the StratL algorithm to address this problem. StratL is based on Learning Sciences and utilizes a student state tracing and tutoring intent selection based on expert-defined transition graphs. We used the advanced but effective learning design Productive Failure in high school math to test it. A field study with high school students demonstrates faithfulness to the learning design and learner-perceived usefulness. We further release a dataset of 20 problems which could serve as a testbed for future improvements.
@@
Scalability. While tutoring can also be explored as a standalone experience, it is important to consider how systems like StratL could be integrated into a classroom or as part of a formal education context. As such, further research should explore (1) how to support teachers in adjusting and personalizing the tutor’s strategy, (2) how to provide relevant feedback to the teacher in order to ensure continuity with classroom content, and (3) how such tutors should be best integrated in existing pedagogical structures.
@@
Social aspects. The social context of a learning experience plays an important role in learning, and “social surround” is even one of the three design layers of Productive Failure (Kapur and Bielaczyc 2012). Generally, supportive and coordinated relationships play an important role in learning (Roschelle and Teasley 1995), and research on how students perceive these tutors at the social level should be explored to paint a complete picture. As such, embodied agents could be considered (Sinha and Malhotra 2022). Beyond cognitive learning outcomes, it is also important to consider affective outcomes of learning activities, such as social-emotional competencies (Alzahrani, Alharbi, and Alodwani 2019), and how these are impacted by the interaction with digital tutors over human ones.
@@
Data Privacy. The field study was approved by the university IRB (Proposal EK-2024-N-256). All data collected were anonymized and no sensitive student information was stored.
6.1 Acknowledgments
We are very grateful to the Math Department of the School of Science and Technology in Singapore for supporting and hosting our field study.
Appendix A Appendix
A.1 Taxonomy and Tutoring Strategy Modeling
@@
We introduce a Learning-Sciences–based taxonomy of tutoring intents denoted $\mathcal{I}$ in Section 3.1. An intent is defined as the goal of a dialog utterance, such as correcting a mistake. The taxonomy groups all possible intents a tutor implemented with StratL can have during the tutoring session. We establish this taxonomy using the theories of Scaffolding (Reiser 2004), Affective Cognition (Lepper et al. 1993) and Productive Failure (Sinha and Kapur 2019; Kapur and Bielaczyc 2012).
@@
Structuring (Reiser 2004) Guide Self-correction Take a second look at your formula.
Correct Are you sure AB is the hypotenuse?
Seek Strategy How would you proceed from now?
Hint Try to use the Pythagorean theorem.
@@
Offload The result of this computation is ….
Problematizing (Reiser 2004) Identify Limits Can you think of a case where it fails?
Prompt Intuition/Hypothesis In your own words, what could be ….
Elicit Articulation Can you formally write this idea?
(Self-)Reflect (Katz, O’Donnell, and Kay 2000) What is the main takeaway of today?
Affective (Lepper et al. 1993; Sinha and Kapur 2019) Maintain Sense of Challenge I dare you to compute ….
@@
Evoke Curiosity In engineering, these equations are used ….
Generic (Macina et al. 2023a) State Greetings/Farewell Hello! How can I help you today?
@@
A.2 Implementation
Student State Tracing
The set $\mathcal{S}$ (as per the notation used in Section 3.2) of labels (also called features) we choose from during the Student State Tracing procedure have been discussed with learning scientists and the prompt structure follows one used by the CLASS framework (Sonkar et al. 2023). We implement the assessment as a multi-label natural language classification task: the last student’s utterance is mapped to a predefined set of features representing the student’s intents, errors’ type, and emotional state. The full list of features can be read from the prompt in Figure 3 (letters a–m in the prompt). We call GPT-4o version ‘gpt-4o-2024-08-06’ on this prompt, with a temperature of 0 and a top_p parameter of $0.1$.
@@ A.2 Implementation
Figure 3: Student State Tracing Prompt to select the student state features, from those enumerated with lowercase letters (a, …, m).
We require the model to generate a justification for its assessment before writing the final selection. This justification serves a double purpose. First, it enhances the system’s explainability. We used these justifications to understand why the LLM selected some erroneous labels and used them to incrementally improve the system, including the formulation of the prompt. Second, this technique usually enhances the accuracy of classification task (Wu, Terry, and Cai 2021). For this purpose, it is important to require the justification to be written before the final label selection, as we use an autoregressive LLM, which can only consider context monodirectionally from the left. In the prompt, we use ‘##’ as a delimiter between the instructions and grounding information (such as the problem and its solution), as suggested in (OpenAi 2024). In lines 2 and 7 we surround “Problem Statement” and “Provided Solution” by ‘*’ to facilitate their reference later in the prompt (Reynolds and McDonell 2021). The core selection prompt from lines 15 to 31 is inspired by the assessment used in the CLASS framework (Sonkar et al. 2023). The exact formulation for each label has been optimized by trial and error over development iterations of the prototype. Lines 33 and 42 summarize the goal and ask the LLM to work step-by-step, a prompting technique drastically improving performance on reasoning benchmarks (Kojima et al. 2022). The last line, containing the first character of the expected answer (‘{’), acts as a syntactical constraint to guide the LLM to follow the JSON format (Reynolds and McDonell 2021).
@@
A.3 Intent Selection
As detailed in Section 3.3 The Intent Selection is the procedure by which the tutor transitions from its previous intents $\mathbf{i}_{n}$ to its next $\mathbf{i}_{n+1}$ conditioned on the student state $\mathbf{s}_{n}$. This transition is represented by a deterministic transition graph. We recall that multiple intents can be selected. To transition from $\mathbf{i}_{n}$ to $\mathbf{i}_{n+1}$. One can follow the following algorithm:
  1. 1.
    For each intent $I$ in $\mathbf{i}_{n}$, start in the corresponding node in the transition graph.
  2. 2.
    The transitions from one node to the other(s) are conditioned on $\mathbf{s}_{n}$. If a transition condition (written on the arrows in the graph) is satisfied by $\mathbf{s}_{n}$, then follow the arrow and add to $\mathbf{i}_{n+1}$ the intent corresponding to the new node reached. Multiple arrows can be followed at the same time, and every corresponding intent is added. If no condition is met, no intent is added ($I$ is not either).
  3. 3.
    Repeat step 2 for each intent $I$ in $\mathbf{i}_{n}$.
For our implementation, we use an Intent Selection defined in collaboration with Productive Failure domain experts. We represent our algorithm by the transition graph in Figure 4. Each node of the graph represents a state, corresponding to an intent of Table 3. Arrows correspond to transitions between states. The letters used in the logical conditions on the arrows correspond to the Student State Tracing features detailed in Figure 3. For example, ‘b’ corresponds to the state feature “The student made an error in the algebraic manipulation”. These labels are combined in logical statements that condition the transitions. Arrows with no starting node represent transitions that can be taken from any state. We represent them this way for the sake of clarity. We recall that we allow for multiple intents, so if multiple transitions are possible at step $t$, they are all followed, and all corresponding intents are considered at step $t+1$. The Learning Science explanations for this graph are detailed in Appendix A.5.
@@ A.3 Intent Selection
Correct Correct the student’s mistake by giving them a hint.
Seek Strategy Encourage and make the student find on their own what is the next step to solve the problem, for example by asking a question.
Hint Give a hint to the student to help them find the next step. Do *not* provide the answer.
@@
Offload Correct and perform the numerical computation for the student.
@@
Prompt Intuition/Hypothesis Ask the student start by providing a guess or explain their intuition of the problem.
Elicit Articulation Ask the student to write their intuition mathematically or detail their answer.
(Self-)Reflect Step back and reflect on the solution. Ask to recapitulate and *briefly* underline more general implications and connections.
@@
Promote Sense of Control Promote a sense of control.
Evoke Curiosity Evoke curiosity.
Generic State Greetings/Farewell Say goodbye and end the conversation
@@
Figure 4: Graph representation of a Productive Failure (Kapur 2008) Intent Selection. How to read this figure is explained in Appendix A.3. The Learning Sciences justifications for this modeling of Productive Failure are given in Appendix A.5. The node labels correspond to the intents in Table 3. The variables used in the arrows’ logical propositions correspond to the state features enumerated with lowercase letters (a, …, m) in Figure 3.
@@
A.4 Intent Dependent Steering
The Intent-Dependend Steering is the final step executed before calling the tutor LLM. It compiles the dialog history, grounding, and the intents selected to generate the prompts for the tutor LLM. In our implementation, these prompts are made of the initial prompt, describing the LLM’s role and persona, of the dialog history, the grounding presenting the exercise and its solution, and of intent-dependent prompt additions aiming at setting the intent of the LLM for its next utterance. The initial prompt is fixed (displayed in Figure 5), and the full dialog history is provided to the LLM. The intent-dependent prompt additions are short messages concatenated at the end of the initial prompt, used to steer the tutor’s intent for the next utterance. We use a one-to-one mapping between the intents and sentences to include in the intent-dependent prompt additions. The precise correspondence between each intent and prompt addition is provided in Table 4.
Line 1 sets the LLM’s persona. Lines 2 to 8 outline general pedagogical guidelines compiled from Learning Sciences research on tutoring (Merrill et al. 1992; Freeman et al. 2014; Lepper et al. 1993). Line 9 is only used to improve the appearance of the tutor’s messages. Line 11 and 19 recapitulate the task for better performance. Lines 13 to 17 provide context about the problem to solve and its solution.

### v2

@@ 5.3 Questionnaire Feedback
Our field test demonstrated the ability of the StratL tutor to follow PF. This section highlights some qualitative insights from the field study feedback.
@@
Robustness As discussed in Section 5.1, the Student State Tracing procedure can make labeling mistakes. These mistakes can then compromise the Intent Selection and result in the selection of contradictory intents. This is exemplified by the feedback of a student: “The AI keeps on asking me to check my answer although it was correct”.
@@
Solution constraint A student assigned to the no intents baseline (V2) complained that the tutor would not let them use a solution entirely different from the expected one: “The AI only listens to key points that it wants, and ignores everything else (e.g. my idea of using […])” A good tutor should encourage original solutions. This highlights a pedagogical problem, as in this case, the tutor is harmful to the learning by restricting the student.
6 Conclusion
@@
We present the pedagogical limitations of LLMs in tutoring as the Pedagogical Steering problem. We introduce StratL, an algorithm addressing this problem. StratL integrates ideas from Learning Sciences with LLMs, utilizing a Student State Tracing and a Tutoring Intent Selection procedure based on expert-defined transition graphs to dynamically steer the LLM to follow a modeled pedagogy. A field study with high school students demonstrates StratL’s ability to follow the Productive Failure learning design.
7 Limitations
@@
Integration. While tutoring can also be explored as a standalone experience, it is important to consider how systems like StratL could be integrated into a classroom or as part of a formal education context. As such, further research should explore (1) how to support teachers in adjusting and personalizing the tutor’s strategy, (2) how to provide relevant feedback to the teacher in order to ensure continuity with classroom content, and (3) how such tutors should be best integrated in existing pedagogical structures.
Risks. As students increasingly utilize AI tools in their learning process, we must ensure they are beneficial to learning. Some studies empirically demonstrated that generative AI tools could harm learning Bastani et al. (2024), through mechanisms such as overreliance Zhai et al. (2024). With our work, we mitigate these effects by empowering teachers and learning scientists to control their student’s interactions with generative AI through the use of StratL. However, further field studies on the reliance behavior of students with a StratLLLM tutor should be conducted before large-scale deployment. Wang et al. (2024b) discusses more broadly the risks associated with the use of LLMs in education.
@@
Social aspects. The social context of a learning experience plays an important role in learning, and “social surround” is even one of the three design layers of PF Kapur and Bielaczyc (2012). Generally, supportive and coordinated relationships play an important role in learning Roschelle and Teasley (1995), and research on how students perceive these tutors at the social level should be explored to paint a complete picture. Beyond cognitive learning outcomes, it is also important to consider affective outcomes of learning activities, such as social-emotional competencies Alzahrani et al. (2019), and how these are impacted by the interaction with digital tutors over human ones.
@@
Data Privacy. The field study was approved by the university IRB. All data collected were anonymized and no sensitive student information was stored.
Appendix A Appendix Evaluation Details
A.1 Dataset
We collected 15 problems from the PF literature  Kapur (2008); Kapur and Bielaczyc (2012); Kapur (2010); Kerrigan et al. (2021); Kerrigan (2018); Kapur and Toh (2015); Loibl and Rummel (2014); Newman and DeCaro (2019); Mazziotti et al. (2016). We present the columns of the dataset in Table 3. We curated three types of problems: Invention problems ask the student to invent a new way to solve a problem. They typically admit several correct but suboptimal answers, as long as the student can justify them. Students are not expected to have previous exposure to the tools canonically used to solve the exercise. The problem Consistency is a good example of invention problem. Ill-structured problems are open-ended problems that contain some non-relevant information and a convoluted story. These problems often admit a single solution, but there can be several ways to reach it. The problem Country is an example of such a problem. Well-structured problems are more direct versions of the ill-structured problems. They resemble classical exercises that students usually solve in a mathematics class. More information about the different types of exercise in PF and their role can be found in Kapur (2008) and Kapur and Bielaczyc (2012). Our dataset is intended to be used as a set of high-quality testing problems for evaluating a tutoring system’s pedagogical skills. We release it under CC-BY-4.0 license [^2] https://creativecommons.org/licenses/by/4.0/deed.en.
A.2 Test Problems
Figures 4 and 5 display the two test problems we used in our field test. They are also included in the dataset we release.
A.3 PF Score and RSM Details
PF Score
We present in the following paragraphs the precise reasoning sub-steps used for the PF Score grading as explained in Section 4.4.
RSM
We tracked the number of student-generated Representation and Solution Methods (RSMs) in problem Consistency. We observed students proposing various solutions to the problem, including comparing the three soccer players based on the variance of their number of goals scored or on their mean absolute year-to-year differences.
Show your work and justify why you think your method is correct.
Solution:
Initial quotas:
- A: 1,646,000 / 50,000 = 32.92 → 32 seats
- B: 6,936,000 / 50,000 = 138.72 → 138 seats
- C: 154,000 / 50,000 = 3.08 → 3 seats
- D: 2,091,000 / 50,000 = 41.82 → 41 seats
- E: 685,000 / 50,000 = 13.70 → 13 seats
- F: 988,000 / 50,000 = 19.76 → 19 seats
Total initial seats assigned = 32 + 138 + 3 + 41 + 13 + 19 = 246
Seats left to distribute = 250 - 246 = 4
Distribute the surplus seats based on largest remainders:
Remainders (from the divisions above):
- A: 0.92
- B: 0.72
- C: 0.08
- D: 0.82
- E: 0.70
- F: 0.76
Figure 4: Problem Country and its solution.
The organizers of the Premier League Federation have to decide which one of the three players Mike Arwen, Dave Backhand and Ivan Right - should receive the "The Most Consistent Player for the Past 5 Years" award. Table 1 shows the number of goals that each striker scored between 2019 and 2023.
(1) Design as many different measures of consistency as you can.
Year Mike Arwen Dave Backhand Ivan Right
2007 13 12 14
2008 12 14 10
2009 15 16 18
2010 17 15 18
2011 13 13 15
Solution:
First, compute the mean:
Mean number of goals for Mike: 14
Mean number of goals for Dave: 14
Mean number of goals for Ivan: 15
Sum squared deviation for Mike: 16
Sum squared deviation for Dave: 10
Sum squared deviation for Ivan: 44
Then divide by the number of data points to get the variance:
Variance for Mike: 12/5 = 3.2
Variance for Dave: 10/5 = 2
Variance for Ivan: 44/5 = 8.8
So according to the variance, Dave is the most consistent player.
Figure 5: Problem Consistency and its solution.
A.4 Student State Tracing Evaluation
$\mu$-average precision $\mu$-average recall $\mu$-average F1
Full version 0.82 0.75 0.77
No Justification 0.71 0.72 0.70
Short Memory 0.64 0.75 0.68
Methodology
We test the procedure using two tutoring conversations of about ten dialog turns for each of the two test problems detailed in Appendix A.2. These conversations are manually annotated to obtain the reference data. As detailed in Appendix B.2, we use an LLM to perform this task in our implementation of StratL. This LLM is prompted with the full conversation history and is asked to output a justification for the labels it selected. We compare this implementation to a version not outputting justifications and a version that only keeps the three last utterances in the prompts’ history to increase the processing speed.
As a multi-label classification task, we compute for each label the precision, recall, and F1 score. We aggregate these metrics across labels using an average weighted by the relative cardinality of each label, known as the micro average ($\mu$-average). We report the results in Table 4.
Results
A.5 Instructions given to the Field Test Participants
Figure 6 displays the instructions and consent form given to the field test participants before entering the study.
Participant ID: <ID>. Please write down and keep this ID for future reference.
Contact project team: <REDACTED>
Data Protection Officer: <REDACTED>
We would like to ask you if you are willing to participate in our research project. Your participation is voluntary. Please read the text below carefully and ask the contact person about anything you do not understand or would like to know.
##### What is investigated and how?
You will have the opportunity to test our AI math tutor. We will investigate whether our AI tutor is able to follow a given pedagogical strategy. You will be asked to solve a math exercise with the help of the tutor. We will analyze your tutoring session to extract some numerical metrics related to how good the tutor was. The results will be aggregated into statistics. You will then give some feedback on the tutor.
##### Who can participate?
You need to be able to write early high school-level mathematics in English to participate.
##### What am I supposed to do as a participant?
As a participant, you will connect to the website provided by your teacher. You will find a chat interface with an AI math tutor on this website. You will first read the AI tutorial, and then chat by text with the AI tutor to solve the exercise. After 15 minutes with the tutor, you will then fill in a feedback form on your experience with the tutor.
##### What are my rights during participation?
Your participation in this study is voluntary. You may withdraw your participation at any time without specifying reasons and without any disadvantages.
##### What risks and benefits can I expect?
We don’t expect any risk for you. Your conversation with the tutor will be recorded in an anonymous fashion and won’t be shared with anyone beyond the researchers involved in the project. The AI tutor might sometimes make mistakes. The tutor is programmed to follow a pedagogy that aims to improve your understanding of math. This session can thus be a very good practice session for math.
##### Will I be compensated for participating?
You will not be compensated for participating in this study. This educational activity will help you learn math.
##### What data is collected from me and how is it used?
Your tutoring session (the messages you send to the AI and the ones you receive) will be recorded. The post-test feedback form will ask you your gender, if you like math, and will ask you to rate your experience with the AI tutor. Only the authors will have access to this data, and it will be anonymous.
Members of the <REDACTED> may access the original data for examination purposes. Strict confidentiality will be observed at any time.
Results will be published in conferences in a completely anonymous fashion. The conversations with the tutor will not be published nor shared beyond the project team. The data will only be published through aggregated statistics and in anonymized form and thus no conclusions can be drawn about individuals.
The anonymized data will be stored on a password-protected encrypted computer and a private, password-protected online repository. All anonymized data will be stored for a maximum of two years after the end of the study.
##### What are my rights to my personal data?
Before the irrevocable anonymization of the collected data, you can request information about the personal data collected from you at any time and without giving reasons. You can also request that it be rectified, handed over to you, barred for processing or erased. To do so, please contact the person indicated above. Please write down and remember your participant ID indicated at the top of this form as it will be required to identify your anonymous data.
The Federal Data Protection Act (FADP), and if applicable the European General Data Protection Regulation (GDPR)) are complied with when collecting and processing personal data.
##### How am I insured?
You are insured for adverse health effects that are directly caused by the study and can be demonstrated to be attributable to fault on the part of the project team or <REDACTED>. You are responsible for insuring yourself against any other adverse health effects such as might occur, for instance, in connection with the trip to or from the place where the study is conducted.
##### Who funds this study? Who reviewed this study?
The study is funded by <REDACTED>.
This study was examined by the <REDACTED>.
The secretariat of the <REDACTED> is available to help you with complaints in connection with your participation. Contact: <REDACTED>
## Consent Form:
I, the participant, confirm by clicking the following button that:
- I have read and understood the study information. My questions have been answered completely and to my satisfaction.
- I comply with the inclusion criteria for participation described above. I am aware of the requirements and restrictions to be observed during the study.
- I have had enough time to decide about my participation.
- I participate in this study voluntarily and consent that my personal data be used as described above.
- I understand that I can stop participating at any moment.
- I have written down my participant ID and understand that it is the only way that the authors can identify my data for any request I can have, as my data is anonymized.
##### Participant ID: <ID>
##### By clicking the button, I confirm that I have read the above information and agree to it.
Figure 6: Instructions and consent form signed by the field test participants
A.6 User interface
Figure 7 displays the chat interface of our prototype tutor.
Appendix B Appendix Learning Sciences and Implementation Details
We release our code under the MIT License in the supplementary material.
B.1 Intents Taxonomy
@@
We introduce a Learning-Sciences–based taxonomy of tutoring intents denoted $\mathcal{I}$ in Section 3.1. An intent is defined as the goal of a dialog utterance, such as correcting a mistake. The taxonomy groups all possible intents a tutor implemented with StratL can have during the tutoring session. We establish this taxonomy using the theories of Scaffolding Reiser (2004), Affective Cognition Lepper et al. (1993) and PF Sinha and Kapur (2019); Kapur and Bielaczyc (2012).
@@
Structuring Reiser (2004) Guide Self-correction Take a second look at your formula.
Correct the Mistake Are you sure AB is the hypotenuse?
Ask for the Next Step How would you proceed from now?
Give a Hint Try to use the Pythagorean theorem.
@@
Perform a Computation The result of this computation is ….
Problematizing Reiser (2004) Identify Limits Can you think of a case where it fails?
Ask for Intuition In your own words, what could be ….
Ask for Articulation Can you formally write this idea?
Guide Self-Reflection Katz et al. (2000) What is the main takeaway of today?
Affective Lepper et al. (1993); Sinha and Kapur (2019) Maintain Sense of Challenge I dare you to compute ….
@@
Evoke Curiosity These equations are also used in ….
Generic Macina et al. (2023a) Say Goodbye Congratulations! See you next time.
@@
B.2 Prompts and Hyperparameters
B.2.1 Student State Tracing
The set $\mathcal{S}$ (as per the notation used in Section 3.1) of labels (also called features) we choose from during the Student State Tracing procedure have been discussed with learning scientists and the prompt structure follows one used by the CLASS framework Sonkar et al. (2023). We implement the assessment as a multi-label natural language classification task: the last student’s utterance is mapped to a predefined set of features representing the student’s intents, errors’ type, and emotional state. The full list of features can be read from the prompt in Figure 8 (letters a–m in the prompt). We call GPT-4o version ‘gpt-4o-2024-08-06’ on this prompt, with a temperature of 0 and a top_p parameter of $0.1$.
@@ B.2.1 Student State Tracing
Figure 8: Student State Tracing Prompt to select the student state features, from those enumerated with lowercase letters (‘a’, …, ‘m’).
We require the model to generate a justification for its assessment before writing the final selection. This justification serves a double purpose. First, it enhances the system’s explainability. We used these justifications to understand why the LLM selected some erroneous labels and used them to incrementally improve the system, including the formulation of the prompt. Second, this technique usually enhances the accuracy of classification task Wu et al. (2021). For this purpose, it is important to require the justification to be written before the final label selection, as we use an autoregressive LLM, which can only consider context monodirectionally from the left. In the prompt, we use ‘##’ as a delimiter between the instructions and grounding information (such as the problem and its solution), as suggested in OpenAi (2024). In lines 2 and 7 we surround “Problem Statement” and “Provided Solution” by ‘*’ to facilitate their reference later in the prompt Reynolds and McDonell (2021). The core selection prompt from lines 15 to 31 is inspired by the assessment used in the CLASS framework Sonkar et al. (2023). The exact formulation for each label has been optimized by trial and error over development iterations of the prototype. Lines 33 and 42 summarize the goal and ask the LLM to work step-by-step, a prompting technique drastically improving performance on reasoning benchmarks Kojima et al. (2022). The last line, containing the first character of the expected answer (‘{’), acts as a syntactical constraint to guide the LLM to follow the JSON format Reynolds and McDonell (2021).
@@
Correct the Mistake Correct the student’s mistake by giving them a hint.
Ask for Intuition Encourage and make the student find on their own what is the next step to solve the problem, for example by asking a question.
Give a Hint Give a hint to the student to help them find the next step. Do *not* provide the answer.
@@
Perform a Computation Correct and perform the numerical computation for the student.
@@
Ask for Intuition Ask the student start by providing a guess or explain their intuition of the problem.
Ask for Articulation Ask the student to write their intuition mathematically or detail their answer.
Guide Self-Reflection Step back and reflect on the solution. Ask to recapitulate and *briefly* underline more general implications and connections.
@@
Promote Sense of Control Make sure your student feels in control.
Evoke Curiosity Promote your student’s curiosity.
Generic Say Goodbye Say goodbye and end the conversation
@@
B.2.2 Intent Dependent Steering
The Intent-Dependend Steering is the final step executed before calling the tutor LLM. It compiles the dialog history, grounding, and the intents selected to generate the prompts for the tutor LLM. The generated prompts are made of the initial prompt, describing the LLM’s role and persona, of the dialog history, the grounding presenting the exercise and its solution, and of intent-dependent prompt additions aiming at setting the intent of the LLM for its next utterance. The initial prompt is fixed (displayed in Figure 9), and the full dialog history is provided to the LLM. The intent-dependent prompt additions are short messages concatenated at the end of the initial prompt, used to steer the tutor’s intent for the next utterance. We use a one-to-one mapping between the intents and sentences to include in the intent-dependent prompt additions. Table 6 displays the intent to prompt addition correspondence.
In Figure 9, line 1 sets the LLM’s persona. Lines 2 to 8 outline general pedagogical guidelines compiled from Learning Sciences research on tutoring Merrill et al. (1992); Freeman et al. (2014); Lepper et al. (1993). Line 9 is only used to improve the appearance of the tutor’s messages. Line 11 and 19 recapitulate the task for better performance. Lines 13 to 17 provide context about the problem to solve and its solution.